# Small-Time Local Controllability of the multi-input bilinear Schrödinger equation thanks to a quadratic term

### v4

@@ 2.6 Tangent vector
One considers $T>0$, and $T_{1}:=\frac{T}{2}$. Then, the previous proposition gives the result. ∎
@@ 2. $(A_{2})$
    Then, for all $x_{0}\in\mathbb{R}^{d}$, $T\mapsto\mathrm{d}\mathscr{F}_{T}(0,(0,0))(x_{0},(0,0))=e^{TDf_{0}(0)}x_{0}$ is continuous on $\mathbb{R}$, and $\mathrm{d}\mathscr{F}_{0}(0,(0,0))(x_{0},(0,0))=x_{0}$.
@@ 4. $(A_{4})$
    By hypothesis, $H=\text{Im}(\mathscr{F}_{T}(0,0)\cdot(0,\cdot))$ is a closed subspace (finite dimension), this is the reachable set of the linearized system around 00. This space doesn’t depend on $T$, and its codimension is finite.
@@ 5. $(A_{5})$
    Finally, Proposition 2.24 shows that, for all $j\in\llbracket r+1,d\rrbracket$, $f_{b_{j}}(0)$ is a small-time $W^{m,\infty}_{0}$-continuously approximately reachable vector. Then, as a supplementary space of $H$ is given by $\text{Span}\left(\left(f_{b_{j}}(0)\right)_{r+1\leqslant j\leqslant d}\right)$, this condition is verified.
@@ 3 Preliminaries on the Schrödinger equation
In all the document, we will note $\omega_{j}:=\lambda_{j}-\lambda_{1}$, and $\nu_{j}:=\lambda_{K}-\lambda_{j}$, with $j\geqslant 1$, $K\geqslant 2$. For $\varphi\in L^{2}(0,1)$, $t\in\mathbb{R}$, one defines
@@ 3.1 Well-posedness
Let $T>0$, $\mu_{1},\mu_{2}$ satisfying $\mathbf{(H)_{reg}}$, $u,v\in H^{m}_{0}((0,T),\mathbb{R})$, $\psi_{0}\in H^{2(p+m)+3}_{(0)}(0,1)$, and $f\in H^{m}_{0}\left((0,T),H^{2p+3}\cap H^{2p+1}_{(0)}(0,1)\right)$. There exists a unique weak solution to the following equation
@@
Let $T>0$, $\mu_{1},\mu_{2}$ satisfying $\mathbf{(H)_{reg}}$, and $\psi_{0}\in H^{2(p+m)+3}_{(0)}(0,1)$. For all $R>0$, there exists $C=C(T,\mu_{1},\mu_{2},R)>0$ such that for all $u,v\in H^{m}_{0}(0,T)$ with $\left\|u\right\|_{H^{m}},\left\|v\right\|_{H^{m}}<R$, one has
$\left\|\psi(T;(u,v),\psi_{0}+\varphi_{1})-\psi(T;(u,v),\varphi_{1})-\psi(T;(0,0),\psi_{0})\right\|_{H_{(0)}^{2(p+m)+3}}\\ \leqslant C\left\|(u,v)\right\|_{H^{m}}\left\|\psi_{0}\right\|_{H_{(0)}^{2(p+m)+3}}.$ (3.5)
@@
When $m\geqslant 1$, the Sobolev space $H^{m}(0,T)$ has an algebra structure, and one obtains the following inequality
@@ 3.2 Expansion of the solution
In the framework of this article, $\mathbf{(H)_{lin,K,1}}$ implies $\langle\Psi(T),\varphi_{K}\rangle\equiv 0$, for every controls $(u,v)$. To determine the evolution of $\psi$ in this direction, we refine the approximation, and extend the development of the solution to the quadratic term. The second-order term $\xi\in\mathcal{C}^{m}\left([0,T],H^{2p+3}_{(0)}(0,1)\right)$ is the solution to the following system
@@
the first term being 00 thanks to orthogonality ($K\neq 1$), and the second one by hypothesis $\mathbf{(H)_{lin,K,1}}$. For $1\leqslant\alpha,\beta\leqslant 2$, one defines
@@
As $\psi_{1}(t)\in H^{2p+3}_{(0)}(0,1)$, and $\mu_{1},\mu_{2}$ verify $\mathbf{(H)_{reg}}$, Remark 3.3 ensures that $(u(t)\mu_{1}+v(t)\mu_{2})\psi_{1}(t)\in H^{2p+3}\cap H^{2p+1}_{(0)}(0,1)$. Then, we can use the regularization inequality given by (3.3) with $m=0$ to obtain
@@
Computing this norm, and using the algebra structure of the Sobolev space $H^{2p+3}(0,1)$,
$\left\|u\mu_{1}\left(\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}\right)\right\|_{L^{2}\left((0,T),H^{2p+3}\right)}\\ \leqslant C\left\|\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}\right\|_{L^{\infty}\left((0,T),H^{2p+3}_{(0)}\right)}\left\|\mu_{1}\right\|_{H^{2p+3}}\left\|u\right\|_{L^{2}(0,T)}.$ (3.17)
@@
$\left\|u\mu_{1}\left(\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}-\Psi\right)\right\|_{L^{2}\left((0,T),H^{2p+3}\right)}\\ \leqslant C\left\|\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}-\Psi\right\|_{L^{\infty}\left((0,T),H^{2p+3}_{(0)}\right)}\left\|\mu_{1}\right\|_{H^{2p+3}}\left\|u\right\|_{L^{2}(0,T)}.$ (3.18)
@@ 3.3 Control in projection
Let $p,m,k\geqslant 0$, $K\geqslant 2$, such that $k\leqslant p$, $\mu_{1},\mu_{2}$ functions satisfying $\mathbf{(H)_{reg}}$, and $\mathbf{(H)_{lin,K,2}}$. Then, the bilinear Schrödinger equation (1.1) is $H^{m}_{0}-$STLC in projection on $\mathcal{H}$ around the ground state in $H^{2(p+m)+3}_{(0)}$. More precisely, for any $T_{1}<T$, there exist $\delta,C>0$, and a $\mathcal{C}^{1}$ map $\Gamma_{T_{1},T}:\mathcal{V}_{T_{1}}\times\mathcal{V}_{T}\to H^{m}_{0}((T_{1},T),\mathbb{R})^{2}$, where
@@
such that, $\Gamma_{T_{1},T}(\psi_{1}(T_{1}),\psi_{1}(T))=(0,0)$ and for every $\psi_{0},\psi_{f}\in\mathcal{V}_{T_{1}}\times\mathcal{V}_{T}$ , the solution to (1.1) on $(T_{1},T)$ with initial data $\psi_{0}$ in $t=T_{1}$, and controls $u,v:=\Gamma_{T_{1},T}(\psi_{0},\psi_{f})$ satisfies
@@ 4 Proof of the main theorem
    In the second step, we correct linearly the system in the other directions $\mathcal{H}$, thanks to Theorem 3.8. Then, we examine the impact of this second step on the correction made in the first one. A priori, the second step can destroy the work of the first, and we use weak estimates to prove that this is not the case.
@@
    We use a trick introduced by Kawski in [35], and executed by Bournissou in [16] to move along the other real direction lost $i^{n+1}\varphi_{K}$.
@@ 4.1 Asymptotic estimates on the quadratic term of the solution
$\left\langle\xi(T),\psi_{K}(T)\right\rangle=-iA_{q+1}^{1}\int_{0}^{T}u_{q+1}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t+i^{n}\gamma_{n}\int_{0}^{T}u_{q+1}(t)v_{n-q}(t)e^{i\omega_{K}t}\mathrm{d}t\\ -iA_{q+1}^{2}\int_{0}^{T}v_{q+1}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t+R_{T,1}^{q+1}(u)+\rho_{T}^{q+1}(u,v)+R_{T,2}^{q+1}(v).$ (4.1)
@@
$\mathcal{F}_{T}^{1}(u)=-u_{1}(T)\sum_{j=1}^{+\infty}c_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}u(s)e^{i\omega_{j}s}\mathrm{d}s\right)+\frac{u_{1}(T)^{2}}{2}e^{i\omega_{K}T}\sum_{j=1}^{+\infty}c_{j}\\ -i\sum_{k=1}^{l+1}A_{k}^{1}\int_{0}^{T}u_{k}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t+R_{T,1}^{l+1}(u).$ (4.2)
@@
Then, with integration by parts, we obtain
@@
Let $T>0$, and assume that $u\in L^{2}((0,T),\mathbb{R})$ such that $u_{i}(T)=0$ for $12\leqslant i\leqslant q+$ 1. If $\mathbf{(H)_{reg}}$ and $\mathbf{(H)_{quad,K,1}}$ hold, then
@@
$\mathcal{F}_{T}^{1}(u)=-u_{1}(T)\sum_{j=1}^{+\infty}c_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}u(s)e^{i\omega_{j}s}\mathrm{d}s\right)+\frac{u_{1}(T)^{2}}{2}e^{i\omega_{K}T}\sum_{j=1}^{+\infty}c_{j}\\ -iA_{q+1}^{1}\int_{0}^{T}u_{q+1}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t+R_{T,1}^{q+1}(u).$
Note that $A_{q+1}^{1}=0$ when $n$ is odd, but not necessarily when $n$ is even. We can manipulate the first term: using integration by parts,
@@
$\mathcal{G}_{T}(u,v)=\sum_{k=1}^{q}(-1)^{k}\gamma_{2k}\mathcal{I}_{T}(u_{k+1},v_{k})+i\sum_{k=0}^{n-q-1}(-1)^{k}\gamma_{2k+1}\mathcal{I}_{T}(u_{k+1},v_{k+1})\\ -u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}v(s)e^{i\omega_{j}s}\mathrm{d}s\right)+\rho_{T}^{q+1}(u,v).$ (4.7)
@@
$A=u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}v(s)e^{i\omega_{j}s}\mathrm{d}s\right)+(-1)^{l+1}\sum_{j=1}^{+\infty}d_{j}\omega_{j}^{l+1}\nu_{j}^{l+1}I^{j}_{T}(u_{l+1},v_{l+1})+\\ \sum_{k=0}^{l}(-1)^{k+1}\sum_{j=1}^{+\infty}d_{j}\omega_{j}^{k}\nu_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k})+i\sum_{k=0}^{l}(-1)^{k+1}\sum_{j=1}^{+\infty}d_{j}\omega_{j}^{k}\nu_{j}^{k+1}\mathcal{I}_{T}(u_{k+1},v_{k+1}).$ (4.8)
@@
$A=u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}v(s)e^{i\omega_{j}s}\mathrm{d}s\right)-\sum_{j=1}^{+\infty}d_{j}\omega_{j}\nu_{j}I^{j}_{T}(u_{1},v_{1})+\\ -\sum_{j=1}^{+\infty}d_{j}\mathcal{I}_{T}(u_{1},v)-i\sum_{j=1}^{+\infty}d_{j}\nu_{j}\mathcal{I}_{T}(u_{1},v_{1}).$
For the induction step, it suffices to do two integration by parts in the formula given by the induction hypothesis. Similarly, we prove that for $0\leqslant l\leqslant q$, for all $u,v\in L^{2}((0,T),\mathbb{R})$ such that $v_{i}(T)=0$ for $11\leqslant i\leqslant l+$ 1,
$B=\sum_{k=0}^{l}(-1)^{k}\sum_{j=1}^{+\infty}\tilde{d_{j}}\omega_{j}^{k}\nu_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k})+i\sum_{k=0}^{l}(-1)^{k}\sum_{j=1}^{+\infty}\tilde{d_{j}}\omega_{j}^{k+1}\nu_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k+1})+\\ (-1)^{l+1}\sum_{j=1}^{+\infty}\tilde{d_{j}}\omega_{j}^{l+1}\nu_{j}^{l+1}I^{j}_{T}(v_{l+1},u_{l+1}).$ (4.9)
@@
$\mathcal{G}_{T}(u,v)=-u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}(-i\omega_{j})^{q+1}\left(\int_{0}^{T}v_{q+1}(s)e^{i\omega_{j}s}\mathrm{d}s\right)+\rho_{T}^{q+1}(u,v)+\\ i(-1)^{q}\gamma_{n}\mathcal{I}_{T}(u_{q+1},v_{q+1}).$
@@
Using the equations (3.14), (4.11), (4.12), and (4.13), we obtain (4.1). ∎

### v5

@@ 2.6 Tangent vector
One considers $T>0$ and $T_{1}:=\frac{T}{2}$. Then, the previous proposition gives the result. ∎
@@ 2. $(A_{2})$
    Then, for all $x_{0}\in\mathbb{R}^{d}$, $T\mapsto\mathrm{d}\mathscr{F}_{T}(0,(0,0))(x_{0},(0,0))=e^{TDf_{0}(0)}x_{0}$ is continuous on $\mathbb{R}$ and $\mathrm{d}\mathscr{F}_{0}(0,(0,0))(x_{0},(0,0))=x_{0}$.
@@ 4. $(A_{4})$
    By hypothesis, $H=\text{Im}(\mathscr{F}_{T}(0,0)\cdot(0,\cdot))$ is a closed subspace (finite dimension), this is the reachable set of the linearized system around 00. This space doesn’t depend on $T$ and its codimension is finite.
@@ 5. $(A_{5})$
    Finally, using (2.9), Proposition 2.24 shows that, for all $j\in\llbracket r+1,d\rrbracket$, $f_{b_{j}}(0)$ is a small-time $W^{m,\infty}_{0}$-continuously approximately reachable vector. Then, as a supplementary space of $H$ is given by $\text{Span}\left(\left(f_{b_{j}}(0)\right)_{r+1\leqslant j\leqslant d}\right)$, this condition is verified.
@@ 3 Preliminaries on the Schrödinger equation
In all the document, we will note $\omega_{j}:=\lambda_{j}-\lambda_{1}$ and $\nu_{j}:=\lambda_{K}-\lambda_{j}$, with $j\geqslant 1$, $K\geqslant 2$. For $\varphi\in L^{2}(0,1)$, $t\in\mathbb{R}$, one defines
@@ 3.1 Well-posedness
Let $T>0$, $\mu_{1},\mu_{2}$ satisfying $\mathbf{(H)_{reg}}$, $u,v\in H^{m}_{0}((0,T),\mathbb{R})$, $\psi_{0}\in H^{2(p+m)+3}_{(0)}(0,1)$ and $f\in H^{m}_{0}\left((0,T),H^{2p+3}\cap H^{2p+1}_{(0)}(0,1)\right)$. There exists a unique weak solution to the following equation
@@
Let $T>0$, $\mu_{1},\mu_{2}$ satisfying $\mathbf{(H)_{reg}}$ and $\psi_{0}\in H^{2(p+m)+3}_{(0)}(0,1)$. For all $R>0$, there exists $C=C(T,\mu_{1},\mu_{2},R)>0$ such that for all $u,v\in H^{m}_{0}(0,T)$ with $\left\|u\right\|_{H^{m}},\left\|v\right\|_{H^{m}}<R$, one has
$\begin{gathered}\left\|\psi(T;(u,v),\psi_{0}+\varphi_{1})-\psi(T;(u,v),\varphi_{1})-\psi(T;(0,0),\psi_{0})\right\|_{H_{(0)}^{2(p+m)+3}}\\ \leqslant C\left\|(u,v)\right\|_{H^{m}}\left\|\psi_{0}\right\|_{H_{(0)}^{2(p+m)+3}}.\end{gathered}$ (3.5)
@@
When $m\geqslant 1$, the Sobolev space $H^{m}(0,T)$ has an algebra structure and one obtains the following inequality
@@ 3.2 Expansion of the solution
In the framework of this article, $\mathbf{(H)_{lin,K,1}}$ implies $\langle\Psi(T),\varphi_{K}\rangle\equiv 0$, for every controls $(u,v)$. To determine the evolution of $\psi$ in this direction, we refine the approximation and extend the development of the solution to the quadratic term. The second-order term $\xi\in\mathcal{C}^{m}\left([0,T],H^{2p+3}_{(0)}(0,1)\right)$ is the solution to the following system
@@
the first term being 00 thanks to orthogonality ($K\neq 1$) and the second one by hypothesis $\mathbf{(H)_{lin,K,1}}$. For $1\leqslant\alpha,\beta\leqslant 2$, one defines
@@
As $\psi_{1}(t)\in H^{2p+3}_{(0)}(0,1)$ and $\mu_{1},\mu_{2}$ verify $\mathbf{(H)_{reg}}$, Remark 3.3 ensures that $(u(t)\mu_{1}+v(t)\mu_{2})\psi_{1}(t)\in H^{2p+3}\cap H^{2p+1}_{(0)}(0,1)$. Then, we can use the regularization inequality given by (3.3) with $m=0$ to obtain
@@
Computing this norm and using the algebra structure of the Sobolev space $H^{2p+3}(0,1)$,
$\begin{gathered}\left\|u\mu_{1}\left(\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}\right)\right\|_{L^{2}\left((0,T),H^{2p+3}\right)}\\ \leqslant C\left\|\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}\right\|_{L^{\infty}\left((0,T),H^{2p+3}_{(0)}\right)}\left\|\mu_{1}\right\|_{H^{2p+3}}\left\|u\right\|_{L^{2}(0,T)}.\end{gathered}$ (3.17)
@@
$\begin{gathered}\left\|u\mu_{1}\left(\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}-\Psi\right)\right\|_{L^{2}\left((0,T),H^{2p+3}\right)}\\ \leqslant C\left\|\psi(\cdot;(u,v),\varphi_{1})-\psi_{1}-\Psi\right\|_{L^{\infty}\left((0,T),H^{2p+3}_{(0)}\right)}\left\|\mu_{1}\right\|_{H^{2p+3}}\left\|u\right\|_{L^{2}(0,T)}.\end{gathered}$ (3.18)
@@ 3.3 Control in projection
Let $p,m,k\geqslant 0$, $K\geqslant 2$, such that $k\leqslant p$, $\mu_{1},\mu_{2}$ functions satisfying $\mathbf{(H)_{reg}}$ and $\mathbf{(H)_{lin,K,2}}$. Then, the bilinear Schrödinger equation (1.1) is $H^{m}_{0}-$STLC in projection on $\mathcal{H}$ around the ground state in $H^{2(p+m)+3}_{(0)}$. More precisely, for any $T_{1}<T$, there exist $\delta,C>0$ and a $\mathcal{C}^{1}$ map $\Gamma_{T_{1},T}:\mathcal{V}_{T_{1}}\times\mathcal{V}_{T}\to H^{m}_{0}((T_{1},T),\mathbb{R})^{2}$, where
@@
such that, $\Gamma_{T_{1},T}(\psi_{1}(T_{1}),\psi_{1}(T))=(0,0)$ and for every $\psi_{0},\psi_{f}\in\mathcal{V}_{T_{1}}\times\mathcal{V}_{T}$ , the solution to (1.1) on $(T_{1},T)$ with initial data $\psi_{0}$ in $t=T_{1}$ and controls $u,v:=\Gamma_{T_{1},T}(\psi_{0},\psi_{f})$ satisfies
@@ 4 Proof of the main theorem
    In the second step, we correct linearly the system in the other directions $\mathcal{H}$, thanks to Theorem 3.8. Then, we examine the impact of this second step on the correction made in the first one. A priori, the second step can destroy the work of the first and we use weak estimates to prove that this is not the case.
@@
    We use a trick introduced by Kawski in [35] and executed by Bournissou in [16] to move along the other real direction lost $i^{n+1}\varphi_{K}$.
@@ 4.1 Asymptotic estimates on the quadratic term of the solution
$\begin{gathered}\left\langle\xi(T),\psi_{K}(T)\right\rangle=-iA_{q+1}^{1}\int_{0}^{T}u_{q+1}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t+i^{n}\gamma_{n}\int_{0}^{T}u_{q+1}(t)v_{n-q}(t)e^{i\omega_{K}t}\mathrm{d}t\\ -iA_{q+1}^{2}\int_{0}^{T}v_{q+1}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t+R_{T,1}^{q+1}(u)+\rho_{T}^{q+1}(u,v)+R_{T,2}^{q+1}(v).\end{gathered}$ (4.1)
@@
$\begin{gathered}\mathcal{F}_{T}^{1}(u)=-u_{1}(T)\sum_{j=1}^{+\infty}c_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}u(s)e^{i\omega_{j}s}\mathrm{d}s\right)+\frac{u_{1}(T)^{2}}{2}e^{i\omega_{K}T}\sum_{j=1}^{+\infty}c_{j}\\ -i\sum_{k=1}^{l+1}A_{k}^{1}\int_{0}^{T}u_{k}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t+R_{T,1}^{l+1}(u).\end{gathered}$ (4.2)
@@
Then, with one integration by parts, we obtain
@@
Let $T>0$ and assume that $u\in L^{2}((0,T),\mathbb{R})$ such that $u_{i}(T)=0$ for $12\leqslant i\leqslant q+$ 1. If $\mathbf{(H)_{reg}}$ and $\mathbf{(H)_{quad,K,1}}$ hold, then
@@
$\begin{gathered}\mathcal{F}_{T}^{1}(u)=-u_{1}(T)\sum_{j=1}^{+\infty}c_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}u(s)e^{i\omega_{j}s}\mathrm{d}s\right)+\frac{u_{1}(T)^{2}}{2}e^{i\omega_{K}T}\sum_{j=1}^{+\infty}c_{j}\\ -iA_{q+1}^{1}\int_{0}^{T}u_{q+1}^{2}(t)e^{i\omega_{K}t}\mathrm{d}t+R_{T,1}^{q+1}(u).\end{gathered}$
Note that $A_{q+1}^{1}=0$ when $n$ is odd, but not necessarily when $n$ is even. We can manipulate the first term: using integrations by parts,
@@
$\begin{gathered}\mathcal{G}_{T}(u,v)=\sum_{k=1}^{q}(-1)^{k}\gamma_{2k}\mathcal{I}_{T}(u_{k+1},v_{k})+i\sum_{k=0}^{n-q-1}(-1)^{k}\gamma_{2k+1}\mathcal{I}_{T}(u_{k+1},v_{k+1})\\ -u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}v(s)e^{i\omega_{j}s}\mathrm{d}s\right)+\rho_{T}^{q+1}(u,v).\end{gathered}$ (4.7)
@@
$\begin{gathered}A=u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}v(s)e^{i\omega_{j}s}\mathrm{d}s\right)+(-1)^{l+1}\sum_{j=1}^{+\infty}d_{j}\omega_{j}^{l+1}\nu_{j}^{l+1}I^{j}_{T}(u_{l+1},v_{l+1})\\ +\sum_{k=0}^{l}(-1)^{k+1}\sum_{j=1}^{+\infty}d_{j}\omega_{j}^{k}\nu_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k})+i\sum_{k=0}^{l}(-1)^{k+1}\sum_{j=1}^{+\infty}d_{j}\omega_{j}^{k}\nu_{j}^{k+1}\mathcal{I}_{T}(u_{k+1},v_{k+1}).\end{gathered}$ (4.8)
@@
$\begin{gathered}A=u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}\left(\int_{0}^{T}v(s)e^{i\omega_{j}s}\mathrm{d}s\right)-\sum_{j=1}^{+\infty}d_{j}\omega_{j}\nu_{j}I^{j}_{T}(u_{1},v_{1})\\ -\sum_{j=1}^{+\infty}d_{j}\mathcal{I}_{T}(u_{1},v)-i\sum_{j=1}^{+\infty}d_{j}\nu_{j}\mathcal{I}_{T}(u_{1},v_{1}).\end{gathered}$
For the induction step, it suffices to do two integrations by parts in the formula given by the induction hypothesis. Similarly, we prove that for $0\leqslant l\leqslant q$, for all $u,v\in L^{2}((0,T),\mathbb{R})$ such that $v_{i}(T)=0$ for $11\leqslant i\leqslant l+$ 1,
$\begin{gathered}B=\sum_{k=0}^{l}(-1)^{k}\sum_{j=1}^{+\infty}\tilde{d_{j}}\omega_{j}^{k}\nu_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k})+i\sum_{k=0}^{l}(-1)^{k}\sum_{j=1}^{+\infty}\tilde{d_{j}}\omega_{j}^{k+1}\nu_{j}^{k}\mathcal{I}_{T}(u_{k+1},v_{k+1})\\ +(-1)^{l+1}\sum_{j=1}^{+\infty}\tilde{d_{j}}\omega_{j}^{l+1}\nu_{j}^{l+1}I^{j}_{T}(v_{l+1},u_{l+1}).\end{gathered}$ (4.9)
@@
$\begin{gathered}\mathcal{G}_{T}(u,v)=-u_{1}(T)\sum_{j=1}^{+\infty}d_{j}e^{i\nu_{j}T}(-i\omega_{j})^{q+1}\left(\int_{0}^{T}v_{q+1}(s)e^{i\omega_{j}s}\mathrm{d}s\right)+\rho_{T}^{q+1}(u,v)\\ +i(-1)^{q}\gamma_{n}\mathcal{I}_{T}(u_{q+1},v_{q+1}).\end{gathered}$
@@
Using the equations (3.14), (4.11), (4.12) and (4.13), we obtain (4.1). ∎